Define uniform convergence of sequence of functions. Give an example


1
Expert's answer
2021-03-30T10:44:16-0400

Define uniform convergence of sequence of functions. Give an example

Defifinition. A sequence of functions "f_n:X\to R" converges uniformly to the function "f:X\to R" if and only if "\lim\limits_{n\to+\infty}\sup\{|f_n(x)-f(x)|:x\in X\}=0" or, equivalently, if and only if for any "\varepsilon>0" there exists a sufficiently large integer N such that for all "x\in X" and "n>N" "|f_n(x)-f(x)|<\varepsilon".


Example 1. X=[0, q], where 0<q<1, "f_n(x)=x^n", "f(x)=0". Then for all "x\in X" we have

"|f_n(x)-f(x)|=x^n\leq q^n\to0", hence "\lim\limits_{n\to+\infty}\sup\{|f_n(x)-f(x)|:x\in X\}=0" and the sequence of functions "f_n(x)" converges to zero uniformly.


Example 2. X=[0, 1), "f_n(x)=x^n", "f(x)=0". Then for all "x\in X" we have "|f_n(x)-f(x)|=x^n", "\sup\{x^n:x\in X\}=1", hence "\lim\limits_{n\to+\infty}\sup\{|f_n(x)-f(x)|:x\in X\}=1" and the sequence of functions "f_n(x)" does not converge to f(x) uniformly.